OpenStudy (anonymous):

d/dx integral sin t^2 from 2 to x

zepdrix (zepdrix):

The `Fundamental Theorem of Calculus, Part 1` states:\[\Large\rm \frac{d}{dx}\int\limits_a^x f(t)~dt ~=~f(x)\]

zepdrix (zepdrix):

So for this problem we recognize that our function of t is,\[\Large\rm f(t)=\sin t^2\]So what do you think your result will be? :o